# Steering Towards Fairness: Mitigating Political Bias in LLMs

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have enabled their widespread use across diverse real-world applications. However, concerns remain about their tendency to encode and reproduce ideological biases—particularly along political and economic dimensions. In this paper, we propose a framework for probing and mitigating such biases in decoder-based LLMs through analysis of internal model representations. Grounded in the Political Compass Test (PCT), our method uses contrastive pairs to extract and compare hidden layer activations from models like Mistral and DeepSeek. We introduce a comprehensive activation extraction pipeline capable of layer-wise analysis across multiple ideological axes, revealing meaningful disparities linked to political framing. Our results show that decoder LLMs systematically encode representational bias across layers, which can be leveraged for effective steering vector-based mitigation. This work provides new insights into how political bias is encoded in LLMs and offers a principled approach to debiasing beyond surface-level output interventions.
@@ 1 Introduction
Large Language Models (LLMs) have become foundational tools across a wide spectrum of applications, yet their outputs frequently reflect political and ideological biases, especially in contexts involving sensitive framing or policy-oriented discourse Zheng et al. (2023). This problem is particularly pressing in multilingual regions like Pakistan, where LLMs often produce uneven or culturally misaligned outputs across different languages, amplifying social or political asymmetries Kumar et al. (2023).
@@
To address this gap, we investigate a modular activation-based mitigation framework that uses contrastive ideological prompts from the PCT to extract, analyze, and intervene on latent bias directions within decoder LLMs. At the core of the method is the use of Steering Vector Ensembles (SVE): layer-specific vector representations that capture ideological framing and allow for inference-time debiasing without fine-tuning. Our contributions are as follows:
@@
  • Our pipeline supports scalable extraction and aggregation of hidden activations across decoder LLMs (e.g., Mistral, DeepSeek) in low-resource Pakistani languages.
@@ 2.1 Bias Evaluation via the Political Compass Test (PCT)
Recent research has highlighted that language plays a key role in shaping LLM bias. Thapa et al. (2023) translated the PCT into Nepali and found that smaller models exhibited economic-right bias, while larger ones leaned socially left. Nadeem et al. (2025) extended this analysis to five Pakistani languages (Urdu and Punjabi), showing that models exhibited stronger authoritarian tendencies when generating in low-resource regional languages. Similarly, Helwe et al. (2025) evaluated 15 multilingual LLMs across 50 countries and found that both prompting language and persona assignment significantly influenced model stance—often more so than the nominal national identity Feng et al. (2023).
@@ 2.2 Steering Vectors and Ensemble Approaches for Mitigation
More recent work introduced Steering Vector Ensembles (SVE) Siddique et al. (2025), which aggregate vectors across multiple demographic groups, model layers, or task settings. These ensembles offer improved robustness and generalizability. However, SVE studies have been narrow in scope, often focusing on: Encoder or encoder-decoder architectures like BERT or T5;, Domain-specific settings, such as toxicity or fairness in QA;, and English-only applications, with little attention to ideological framing or multilingual dynamics. Thus, the potential of SVE for open-ended political discourse, especially in decoder LLMs, remains largely unexplored.
@@
  • Lack of systematic contrastive activation pipelines: There is no open-source or standardized pipeline for extracting contrastive activations (e.g., liberal vs. authoritarian) across layers and prompts in decoder LLMs, especially for multilingual bias detection.
  • Underutilization of SVE in political contexts: Steering Vector Ensembles have shown promise in fairness-related domains, but their application to ideological bias mitigation, especially across languages and political axes, remains under-investigated Chen et al. (2020).
To address these limitations, our work introduces an activation-based bias mitigation pipeline tailored for decoder LLMs. Our core contributions include: A scalable, multilingual framework for layer-wise activation extraction using PCT-based contrastive pairs; A representation-level analysis method that identifies and aggregates ideological bias directions; And the first integration of Steering Vector Ensembles into decoder-based LLMs for mitigating political bias across both social and economic axes. By bridging output-based evaluation and internal mitigation strategies, we provide a new foundation for probing and correcting ideological bias in multilingual generative models.
@@ 3.2 Multilingual PCT Dataset Preparation
We build on the multilingual PCT dataset proposed by Nadeem et al. (2025), which adapts the 62 standard Political Compass Test (PCT) statements into five low-resource Pakistani languages: Urdu and Punjabi Smith et al. (2022). We extend this dataset to include English, resulting in six total languages spanning multiple language families Mostefa et al. (2012).
@@ 3.3 Target Model and Layer Selection
We selected deepseek-llm-7b-chat [^1] https://huggingface.co/deepseek-ai/deepseek-vl-7b-chat, Mistral model due to its strong multilingual capabilities and transparent architecture. We also evaluate Mistral-7B-v0.1222https://huggingface.co/mistralai/Mistral-7B-v0.1 to compare bias behavior across model families.
@@ 3.7 Response Quality Metrics
Model Econ. (Before) Soc. (Before) Econ. (After) Soc. (After)
Mistral-7B-Instruct-v0.2 2.5 1.23 0.0 0.5
DeepSeek-Chat -1.0 -1.23 0.0 0.2
Table 1: Bias scores before and after mitigation across models and ideological axes.
@@ 4.1 Hyperparameter Configuration
Figure 3: Bias reduction performance of Steering Vector Ensembles (SVE) and Individual Steering Vectors (ISV) across Social and Economic dimensions under varying input bias intensities.
@@ 5.1 High Resource Language
In our bias mitigation framework, English serves as the high-resource baseline language due to its extensive training data, well-established benchmarks, and consistent performance across models. We use English to construct contrastive political prompts, calibrate steering vectors, and evaluate baseline bias levels before extending the methodology to low-resource Pakistani languages.
@@
Figure 5 illustrates how bias mitigation effectiveness varies across model layers for both Steering Vector Ensembles (SVE) and Individual Steering Vectors (ISV). SVE for social bias stands out, consistently achieving  50% reduction across all layers and operating in the high-effectiveness zone. ISV for economic bias peaks at layer 16 with  30% reduction but declines. In contrast, SVE for economic bias and ISV for social bias remain below 5% and show little variation. The background shading highlights zones of high (green), moderate (yellow), and low (red) effectiveness, clearly emphasizing the stability and superiority of SVE for mitigating social bias. Figure 7 shows that SVE excels at reducing social bias in DeepSeek, notably for Traditional Values and Immigration, while ISV is more effective on economic prompts like Taxation. The contrast reveals each method’s domain-specific strength, highlighting the benefit of combining them for comprehensive bias mitigation.
@@
Figure 6 illustrates the relationship between steering strength and bias reduction for SVE and ISV methods. A clear optimal point emerges at a steering strength of 1.0, where SVE Social achieves peak effectiveness (50%) and ISV Economic also reaches its maximum (28%). Beyond this threshold, performance gradually declines, indicating that excessive steering may over-correct or destabilize model outputs. SVE Economic, by contrast, exhibits only minimal bias reduction across all strength levels. These results underscore the critical role of hyperparameter tuning—particularly steering strength—in maximizing the effectiveness of vector-based debiasing strategies.
@@
Figure 4 presents a comprehensive comparison of SVE and ISV across key aspects of bias mitigation. SVE for social bias demonstrates steady and effective improvement, achieving up to 50% bias reduction early in the optimization process. It also consistently preserves response quality, maintaining fluency and coherence throughout. In contrast, ISV—particularly for economic bias—shows less stable trends and struggles to match SVE in both fairness and quality. SVE further exhibits adaptability by dynamically leveraging different model layers, especially mid-layer regions, to optimize its steering effect. Additionally, it delivers strong bias reduction with relatively low computational overhead, making it more cost-efficient than ISV, which requires more resources for smaller gains. These results underscore SVE’s advantages in robustness, adaptability, and efficiency. Complementing this, Figure 8 illustrates that ideological distinctions are most pronounced in mid-level layers, aligning with where SVE applies its interventions to guide the model toward more neutral and balanced responses.
Figure 7: Bias reduction performance of Steering Vector Ensemble (SVE) and Individual Steering Vector (ISV) across different constrastive pair types.
The results show in a Table  1 a clear improvement in how both models handle ideological bias after mitigation. Before applying our method, Mistral-7B-Instruct-v0.2 leaned heavily in both economic and social directions, with bias scores as high as 2.5 and 1.23. After mitigation, those scores dropped significantly—closer to neutral—showing that the intervention helped balance its responses. Similarly, DeepSeek-Chat started with a noticeable bias in the opposite direction, but also moved toward neutrality after mitigation. Overall, these changes suggest that the approach is effective in steering both models away from extreme positions and helping them generate more balanced, fair outputs. In the Figure 9 how different models respond to bias mitigation methods. In DeepSeek, using Steering Vector Ensembles (SVE) noticeably improves the quality of responses, especially for social topics in both Urdu and Punjabi. The model becomes more fluent and balanced without losing clarity. In some cases—like Punjabi economic prompts—SVE even lowers the quality. This highlights that not all models benefit from the same debiasing strategy, and choosing the right method depends on the model and language involved.
@@
The Steering Vector Ensembles (SVE) are more effective than Individual Steering Vectors (ISV) for mitigating political bias in decoder-based LLMs. SVE achieved up to 60% bias reduction on socially framed prompts while preserving high response quality. ISV showed moderate success on economic prompts (e.g., 30% on Economic Policy) but was largely ineffective on social ones. Layer-wise analysis revealed that mid-level layers (especially 16 and 20) are most indicative of ideological bias. SVE leverages this by aggregating steering vectors across multiple layers using quality-based weighting, making it more robust and generalizable. Both methods performed best at a steering strength of 1.0. DeepSeek-Chat responded particularly well to SVE, producing neutral and fluent outputs across Urdu and Punjabi. Mistral-7B showed slightly better alignment with ISV on economic axes but suffered quality drops with SVE. Overall, SVE offers a scalable, interpretable, and language-agnostic solution for fair stance generation.
@@ 6 Conclusion
This study presents a practical and effective approach to reducing political bias in language models using contrastive prompts from the Political Compass Test. We show that Steering Vector Ensembles (SVE) work better than Individual Steering Vectors (ISV), especially for socially framed prompts, and do so without hurting response quality. By focusing on mid-layer activations and adjusting steering strength, the method stays both adaptable and efficient. It performs well across different languages and model types, making it suitable for real-world, multilingual use. Overall, our findings offer a solid foundation for building fairer and more balanced language models.
7 Limitations
Despite its effectiveness, our approach has several limitations. First, reliance on fixed Political Compass Test (PCT) statements constrains generalizability to other domains of bias. The steering strength parameter $\alpha$ is manually tuned, which may limit adaptability across models or tasks. Our method modifies only the last-token activation, which may not propagate sufficiently to steer longer generations. Finally, the keyword-based bias detection metric may overlook subtle or implicit ideological framing, especially in multilingual settings. Future work may explore dynamic steering, discourse-level evaluation, and expansion to other bias dimensions such as gender, race, and disability.
8 Ethical Considerations

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have enabled their widespread use across diverse real-world applications. However, concerns remain about their tendency to encode and reproduce ideological biases along political and economic dimensions. In this paper, we employ a framework for probing and mitigating such biases in decoder-based LLMs through analysis of internal model representations. Grounded in the Political Compass Test (PCT), this method uses contrastive pairs to extract and compare hidden layer activations from models like Mistral and DeepSeek. We introduce a comprehensive activation extraction pipeline capable of layer-wise analysis across multiple ideological axes, revealing meaningful disparities linked to political framing. Our results show that decoder LLMs systematically encode representational bias across layers, which can be leveraged for effective steering vector-based mitigation. This work provides new insights into how political bias is encoded in LLMs and offers a principled approach to debiasing beyond surface-level output interventions.
@@ 1 Introduction
Large Language Models (LLMs) have become foundational tools across a wide spectrum of applications, yet their outputs frequently reflect political and ideological biases, particularly in contexts involving sensitive framing or policy-oriented discourse Zheng et al. (2023); Afzoon et al. (2025). This problem is particularly pressing in multilingual low-resource settings, where LLMs often produce uneven or culturally misaligned outputs across different languages, amplifying social or political asymmetries Kumar et al. (2023); Maskey et al. (2025).
@@
To address this gap, we investigate a modular activation-based mitigation framework that uses contrastive ideological prompts from the PCT to extract, analyze, and intervene on latent bias directions within decoder LLMs. At the core of the method is the use of Steering Vector Ensembles (SVE): layer-specific vector representations that capture ideological framing and allow for inference-time debiasing without fine-tuning Siddique et al. (2025). Our contributions are as follows:
@@
  • Our pipeline supports scalable extraction and aggregation of hidden activations across decoder LLMs (e.g., Mistral, DeepSeek) in low-resource languages.
@@ 2.1 Bias Evaluation via the Political Compass Test (PCT)
Recent research has highlighted that language plays a key role in shaping LLM bias. Thapa et al. (2023) translated the PCT into Nepali and found that smaller models exhibited economic-right bias, while larger ones leaned socially left. Nadeem et al. (2025) extended this analysis to low-resource languages (Urdu and Punjabi), showing that models exhibited stronger authoritarian tendencies when generating in low-resource regional languages. Similarly, Helwe et al. (2025) evaluated 15 multilingual LLMs across 50 countries and found that both prompting language and persona assignment significantly influenced model stance—often more so than the nominal national identity Feng et al. (2023).
@@ 2.2 Steering Vectors and Ensemble Approaches for Mitigation
More recent work introduced Steering Vector Ensembles (SVE) Siddique et al. (2025), which aggregate vectors across multiple demographic groups, model layers, or task settings. These ensembles offer improved robustness and generalizability. However, SVE studies have been narrow in scope, often focusing on: Encoder or encoder-decoder architectures like BERT or T5;, Domain-specific settings, such as toxicity or fairness in QA;, and English-only applications, with little attention to ideological framing or multilingual dynamics. Thus, the potential of SVE for open-ended political discourse, particularly in decoder LLMs, remains largely unexplored.
@@
  • Lack of systematic contrastive activation pipelines: There is no open-source or standardized pipeline for extracting contrastive activations (e.g., liberal vs. authoritarian) across layers and prompts in decoder LLMs, particularly for multilingual bias detection.
  • Underutilization of SVE in political contexts: Steering Vector Ensembles have shown promise in fairness-related domains, but their application to ideological bias mitigation, particularly across languages and political axes, remains under-investigated Chen et al. (2020).
To address these limitations, our work introduces an activation-based bias mitigation pipeline tailored for decoder LLMs. Our core contributions include: A scalable, multilingual framework for layer-wise activation extraction using PCT-based contrastive pairs; A representation-level analysis method that identifies and aggregates ideological bias directions. And the first integration of Steering Vector Ensembles into decoder-based LLMs for mitigating political bias across both social and economic axes. By bridging output-based evaluation and internal mitigation strategies, we provide a new foundation for probing and correcting ideological bias in multilingual generative models. Although we focus on Political Compass Test (PCT) prompts, our pipeline is modular and could be extended to other domains such as healthcare or education. The ensemble design also improves robustness to prompt framing by aggregating across multiple paraphrases.
@@ 3.2 Multilingual PCT Dataset Preparation
We build on the multilingual PCT dataset proposed by Nadeem et al. (2025), which adapts the 62 standard Political Compass Test (PCT) statements into low-resource languages: Urdu and Punjabi Smith et al. (2022). We extend this dataset to include English, resulting in six total languages spanning multiple language families Mostefa et al. (2012).
@@ 3.3 Target Model and Layer Selection
We selected deepseek-llm-7b-chat [^1] https://huggingface.co/deepseek-ai/deepseek-vl-7b-chat, Mistral model due to its strong multilingual capabilities and transparent architecture. We also evaluate Mistral-7B-v0.1222https://huggingface.co/mistralai/Mistral-7B-v0.1 to compare bias behavior across model families. We selected layers 8, 12, 16, 20, 24, as layer-wise profiling indicated that mid-level layers encode the strongest ideological signals, whereas early layers capture lexical patterns and very late layers primarily influence fluency.
@@ 4.1 Hyperparameter Configuration
Figure 4: Bias reduction performance of SVE and ISV across Social and Economic dimensions under varying input bias intensities.
@@ 5.1 High Resource Language
In our bias mitigation framework, English serves as the high-resource baseline language due to its extensive training data, well-established benchmarks, and consistent performance across models. We use English to construct contrastive political prompts, calibrate steering vectors, and evaluate baseline bias levels before extending the methodology to low-resource languages.
@@
Figure 3 illustrates how bias mitigation effectiveness varies across model layers for both SVE and ISV. SVE for social bias stands out, consistently achieving  50% reduction across all layers and operating in the high-effectiveness zone. ISV for economic bias peaks at layer 16 with  30% reduction but declines. In contrast, SVE for economic bias and ISV for social bias remain below 5% and show little variation. The background shading highlights zones of high (green), moderate (yellow), and low (red) effectiveness, clearly emphasizing the stability and superiority of SVE for mitigating social bias. Figure 4 shows that SVE consistently reduces social bias, while ISV is relatively stronger on economic prompts, highlighting their complementary roles in bias mitigation.
@@
Figure 5 illustrates the relationship between steering strength and bias reduction for SVE and ISV methods. A clear optimal point emerges at a steering strength of 1.0, where SVE Social achieves peak effectiveness (50%) and ISV Economic also reaches its maximum (28%). Beyond this threshold, performance gradually declines, indicating that excessive steering may over-correct or destabilize model outputs. SVE Economic, by contrast, exhibits only minimal bias reduction across all strength levels. These results underscore the critical role of hyperparameter tuning—particularly steering strength—in maximizing the effectiveness of vector-based debiasing strategies. Figure 7 shows that SVE excels at reducing social bias in DeepSeek, notably for Traditional Values and Immigration, while ISV is more effective on economic prompts like Taxation. The contrast reveals each method’s domain-specific strength, highlighting the benefit of combining them for comprehensive bias mitigation.
Model Econ. (Before) Soc. (Before) Econ. (After) Soc. (After)
Mistral-7B-Instruct-v0.2 2.5 1.23 0.0 0.5
DeepSeek-Chat -1.0 -1.23 0.0 0.2
Table 1: Bias scores before and after mitigation across models and ideological axes on Urdu language.
@@
Figure 6 presents a comprehensive comparison of SVE and ISV across key aspects of bias mitigation. SVE for social bias demonstrates steady and effective improvement, achieving up to 50% bias reduction early in the optimization process. It also consistently preserves response quality, maintaining fluency and coherence throughout. In contrast, ISV—particularly for economic bias—shows less stable trends and struggles to match SVE in both fairness and quality. SVE further exhibits adaptability by dynamically leveraging different model layers, particularly mid-layer regions, to optimize its steering effect. Additionally, it delivers strong bias reduction with relatively low computational overhead, making it more cost-efficient than ISV, which requires more resources for smaller gains. These results underscore SVE’s advantages in robustness, adaptability, and efficiency. Complementing this, Figure 8 illustrates that ideological distinctions are most pronounced in mid-level layers, aligning with where SVE applies its interventions to guide the model toward more neutral and balanced responses.
Figure 7: Bias reduction performance of SVE and ISV across different contrastive pair types.
The results show in a Table  1 a clear improvement in how both models handle ideological bias after mitigation. Before applying our method, Mistral-7B-Instruct-v0.2 leaned heavily in both economic and social directions, with bias scores as high as 2.5 and 1.23. After mitigation, those scores dropped significantly—closer to neutral—showing that the intervention helped balance its responses. Similarly, DeepSeek-Chat started with a noticeable bias in the opposite direction, but also moved toward neutrality after mitigation. Overall, these changes suggest that the approach is effective in steering both models away from extreme positions and helping them generate more balanced, fair outputs. In the Figure 9 how different models respond to bias mitigation methods. In DeepSeek, using SVE noticeably improves the quality of responses, for social topics in both Urdu (Ur) and Punjabi (Pu). The model becomes more fluent and balanced without losing clarity. In some cases like Punjabi economic prompts, SVE lowers the quality. This highlights that not all models benefit from the same debiasing strategy, and choosing the right method depends on the model and language involved.
@@
Figure 9: Bias reduction performance of ISV and SVE on model representations across economic and social axes for DeepSeek and Mistral.
The SVE is effective than ISV for mitigating political bias in decoder-based LLMs. SVE achieved up to 60% reduction on socially framed prompts while preserving response quality, whereas ISV showed moderate gains on economic prompts but was largely ineffective socially. An ablation of ISV, cosine filtering, and ensemble weighting indicates ensembles drive most bias reduction, with ISVs adding targeted improvements on economic prompts. Layer-wise analysis showed mid-level layers carry the strongest ideological signals, and SVE’s quality-weighted aggregation across these layers improved robustness and generalization. Both methods performed best at steering strength. DeepSeek benefited most from SVE, producing neutral, fluent responses across Urdu and Punjabi, while Mistral-7B aligned slightly better with ISV on economic axes but lost quality with SVE. We observed degraded quality in Punjabi economic prompts, likely due to vocabulary sparsity, suggesting the value of language-specific calibration in low-resource settings.
@@ 6 Conclusion
This study proposes a practical method to reduce political bias in LMs using contrastive prompts from the Political Compass Test. SVE outperforms compared to ISV, particularly on socially framed prompts, while preserving response quality. By targeting mid-layer activations with adjustable steering strength, the approach remains efficient and adaptable across models and languages, providing a foundation for fairer multilingual language models.
Limitations
Our approach has several limitations. First, reliance on Political Compass Test (PCT) statements constrains generalizability; although the pipeline is modular and can be applied to other domains (e.g., healthcare, education, gender, race), this remains to be tested. Second, the steering strength parameter ($\alpha$) and layer selection were manually tuned, limiting adaptability across models; automated calibration could improve robustness. Third, modifying only the last-token activation may not sufficiently propagate steering in longer generations, suggesting a need for dynamic or dialogue-aware steering. Fourth, evaluation relies on keyword-based lexicons, which may miss subtle discursive bias; while stance classification was included, human and discourse-level evaluations are needed. Finally, challenges arose in low-resource settings; Punjabi economic prompts showed reduced quality due to sparse vocabulary, and some entanglement between social and economic axes was observed. Future work could explore language-specific calibration, multi-axis steering, and stance-conditional methods to balance neutrality with context-appropriate stances.
Ethical Considerations